# Multimodal AI predicts clinical outcomes of drug combinations from preclinical data

### v1

@@ Discussion
Figure 6: Madrigal predicts effective drug combinations tailored to individual patients. a, Drug combination datasets to which Madrigal is applied, spanning multiple stages of the drug development pipeline, including preclinical studies, ex vivo studies, clinical trials, and post-marketing surveillance. b, Predicting personalized drug combination efficacy using Madrigal’s encoder. The encoder parameters are frozen, and drug encodings, along with dimension-reduced patient-specific features (genomic and/or clinical attributes), are input into a predictor module to generate efficacy predictions. c, Performance of Madrigal in predicting synergistic drug combinations in BeatAML [39], where prediction target is combination synergy (Methods Sec. 5.6). Model evaluation is conducted on randomly held-out patients. Patient-centric and drug-centric denote two ways of calculating AUROC to evaluate the model (Methods Sec. 5.6). Error bars show 95% confidence interval. d, Performance of Madrigal in predicting drug combination efficacy in PDX Encyclopedia[41] under a leave-one-drug-combination-out setup. The prediction target is treatment response (BestAvgResponse), and the evaluation metric is Kendall’s tau correlation. e, Kaplan-Meier survival estimates stratified by predicted treatment response for the (BKM120 + encorafenib) combination. The predictor is trained on other drug combinations, with treatment response as the prediction target (same predictor as in (d)). Two-sided log-rank test; $p$-value $=0.001$. f, Same as (e) but for the (encorafenib + binimetinib) combination. Two-sided log-rank test; $p$-value $<1e\mbox{-}4$. g, Predicted progression-free survival (PFS, TimeToDouble) for individual patient models treated with the (encorafenib + binimetinib) combination. The predictor is trained on other drug combinations, with PFS as the prediction target. Predictions are color-coded by the observed best response category (calculated from response according to mRECIST [41]) of each patient model, SD+ = stable disease or partial response or complete response, PD = progressive disease. h, Same as (g) but for the (BKM120 + encorafenib) combination.
Figure 7: Flexible descriptions of safety outcomes through a LLM decoding module. a, The Madrigal-LLM architecture involves processing the two drugs with modality-specific encoders and the fusion module as Madrigal, plus a pre-trained LLM (Mistral-7B [126]) to obtain encodings for free-text descriptions of safety outcomes. The parameters of the LLM are frozen. The encodings of the drugs and the free-text descriptions are fed into a trainable self-attention layer, and a trainable MLP layer is then applied to concatenated encodings to obtain the final prediction for the safety outcome described in free text. b, Free-text descriptions of safety outcomes are augmented using GPT-4 to obtain more diverse and potentially informative descriptions. We construct a proper prompt and query GPT-4 to generate augmentations as if it were an expert Doctor, Pharmacologist, or Toxicologist (Methods Sec. 4.1). c, Test performance of Madrigal-LLM is shown across unseen safety outcomes for the cases where the model is trained on only the original text descriptions and where it is trained on augmented descriptions. Only original text descriptions of unseen safety outcomes are used to evaluate performance in both cases to ensure a fair comparison. d, Test performance of Madrigal-LLM trained with augmented descriptions across unseen safety outcomes, with different text description types (original and augmented) used for evaluation. The performances are compared against a naive baseline without utilizing text information (Methods Sec. 4). e, Concordance of Madrigal-LLM trained on only training data (same as (c-d)) with a Madrigal model trained on all data, as measured by Spearman’s rank correlation for unseen safety outcomes across the different text descriptions types. Only those safety outcomes and description types that are significantly correlated ($p$-value $<$ 0.05) are shown. Data in the methemoglobinemia (first) row are clamped at x = 0 for ease of visualization. f, Using Madrigal-LLM for realistic query of drug pairs and an outcome description related to anemia. The results are summarized in the rightmost panel.
@@ 1.1 Drug Combination Safety Dataset
We collect datasets from TWOSIDES (2019-11) [45] and DrugBank (2023-01-04) [46]. TWOSIDES is a database derived from the FDA Adverse Event Reporting System (FAERS). FAERS is a comprehensive repository of adverse event and medication error reports submitted to the FDA. To ensure the reliability and relevance of our data, we adhere to widely accepted criteria in the existing literature on drug safety data mining [127, 128]. These criteria include: (1) a minimum of three reports for the pair of drugs that report the side effect; (2) a proportional reporting ratio of at least 2; (3) a mean reporting frequency of 0.01 or higher; and (4) a chi-square value of 3.841 or higher (p-value $\leq 0.05$). Applying these criteria and filtering out safety outcomes with less than 100 samples [50], we have compiled a total of 4,656,138 samples, which include 1,457 unique drugs and 795 unique safety outcomes.
@@ 1.7 MASH Combination Therapies
1.8 Cancer Drugs and PARPi Combination Therapies Dataset
We collect PARPi’s that are FDA-approved or currently under advanced stages of clinical trials that are sufficiently represented (with at least five different other drugs) for at least two outcomes in our dataset. Three PARPi’s are collected of this kind, namely olaparib, talazoparib, and rucaparib, which are associated with 11, 6, and 9 outcomes, respectively.
We then filter curated approved or clinically investigated PARPi combinations in [78, 73, 64] so that both drugs in each combination are present in our data. This filters out non-small-molecule drugs, leaving us with nine combinations for olaparib, 1 combination for talazoparib, and 1 combination for rucaparib.
To curate a list of drugs indicated for cancer, we consider cancer drugs collected by [135] (version 07/02/2024), whose source data comes from the NCI, FDA, EMA, and others. According to the authors, “The intention is to provide researchers, clinicians, and regulators with an easily filtered database of licensed drugs used to treat cancer. Drugs used in cancer treatments to alleviate symptoms or other supportive care uses or used for diagnostic purposes are not included. Investigational agents and experimental treatments used in clinical trials are also not included.” This list is supplemented by two other sources (excluding PARPi’s):
  • Drugs whose ATC codes start with L. As drugs falling under the L02, L03, and L04 classification might have non-cancer indications, we further ensure that these L02, L03, or L04 drugs have cancer-related descriptions in their description, mechanism of action, or indication in DrugBank [46] or DepMap [17].
  • Drugs indicated for cancer as curated in DrugCentral [89], obtained through PrimeKG [90].
We manually examined all drugs that are not approved in the US. We ensured that they are currently approved in Europe, Japan, South Korea, China, Australia, or Germany or are in an advanced stage of clinical trials (at least phase 3) and not discontinued due to safety or efficacy and not with a withdrawal history, using DrugBank and Synapse [136]. In total, we collect 288 cancer drugs of interest. For all drugs, the categorical mechanism of action annotations are extracted from DepMap [17] and CMap [16], and the target annotations are extracted from DrugBank. The specific drug names and their DrugBank IDs are provided in Supplementary Table S2.
The clinically-investigated PARPi combinations are curated from [78].
@@ 1.10 Patient-derived Xenograft Drug Combination Dataset
The efficacy measures we use are TimeToDouble, which corresponds to progression-free survival (time until tumor volume reaches 200% of baseline), and BestAvgResponse, which corresponds to response (minimum value of the average of $\Delta{Vol}_{t}$ from t $=$ 0 to T, for T $\geq$ 10 d). The drugs in the dataset are mapped to DrugBank ID by name and through manual confirmation with literature.
@@ 2.2.1 Initializing and adapting individual modality-specific encoders
    $(\displaystyle L_{\mathsf{BCE}}^{\mathsf{PKG}}=\frac{1}{n_{\mathsf{PKG}}}\Big{(}$$-\sum_{(s,t,r)\in E}\log h^{\mathsf{PKG}}(f^{\mathsf{PKG}}(s,G),f^{\mathsf{PKG}}(t,G),r)$
@@ 2.2.2 Modality alignment with multimodal contrastive learning
In this stage, our objective is to align the representations generated for the same drug from different modalities with the structure modality with a multimodal contrastive representation learning objective. We adopted the InfoNCE objective [143] with minor modifications similar as in [144], and jointly learn all encoders $f^{m},m\in\mathcal{M}$, initialized from stage 1, s.t. the loss
$L_{\mathsf{cont}}=\sum_{m\neq\mathsf{struc.}}L\left(m,\mathsf{struc.}\right)=\sum_{m\neq\mathsf{struc.}}\left(\ell\left(m,\mathsf{struc.}\right)+\ell\left(\mathsf{struc.},m\right)\right),$
where
$\ell\left(m_{u},m_{v}\right)=-\sum_{i=1}^{B}\log\frac{\mathrm{sim}_{(m_{u},m_{v})}(d_{i},d_{i})}{\sum_{j=1}^{B}\left(\mathrm{sim}_{(m_{u},m_{v})}(d_{i},d_{j})+\mathbf{1}_{j\neq i}\cdot\mathrm{sim}_{(m_{u},m_{u})}(d_{i},d_{j})\right)},$
and
$\mathrm{sim}_{(m_{u},m_{v})}(d_{i},d_{j})=\mathrm{exp}(f^{m_{u}}(d_{i})\cdot f^{m_{v}}(d_{j})/\tau),$
is minimized.
@@ 2.2.3 Additional note on modality alignment
Concretely, let $z_{\mathsf{struc.}},z_{\mathsf{PKG}},z_{\mathsf{CV}},z_{\mathsf{trans.}}$ be modality-specific representations encoded by respective encoders $f^{m}$, where $m=\mathsf{struc.},\mathsf{PKG},\mathsf{CV},\mathsf{trans.}$, as defined before. A careful reader might note that the structure modality is anchor-like in the contrastive objective above. Minimizing the sum of InfoNCE objectives between the structure modality and other modalities, respectively, can be viewed as maximizing a lower bound estimate of the sum of mutual information shared between representations of the structure modality and other modalities, respectively [145], i.e.,
$\max_{\{f^{m}\}_{m\in\mathcal{M}}}{\sum_{m\neq\mathsf{struc.}}{I(z_{\mathsf{struc.}};z_{m})}},$
where $I$ denotes mutual information, which directly aims to align the structure modality with all other modalities. Under the assumption of conditional independence structure between other modalities given the structure modality, i.e. $z_{\mathsf{CV}}\perp z_{\mathsf{PKG}}\ |\ z_{\mathsf{struc.}},z_{\mathsf{trans.}}\perp z_{\mathsf{PKG}}\ |\ z_{\mathsf{struc.}},z_{\mathsf{trans.}}\perp z_{\mathsf{CV}}\ |\ z_{\mathsf{struc.}}$, the pairwise mutual information objective is equivalent to:
$\max_{\{f^{m}\}_{m\in\mathcal{M}}}{\sum_{m}{H(z_{m})}-H(z_{\mathsf{struc.}},z_{\mathsf{PKG}},z_{\mathsf{CV}},z_{\mathsf{trans.}})},$
where $H$ denotes entropy, it can thus be interpreted that the maximization of $\sum_{m}{H(z_{m})}$ ensures that each modality retains its inherent variability and richness, and the minimization of $H(z_{\mathsf{struc.}},z_{\mathsf{PKG}},z_{\mathsf{CV}},z_{\mathsf{trans.}})$ as ensuring that the joint representation is compact and has lower redundancy.
@@ 2.2.4 Model finetuning
Given the impressive performance of attention-based fusion in other multimodal learning contexts, particularly in vision-language models [146, 147, 44], and their flexibility of inputs, we adopt a Transformer encoder architecture for modality fusion to model the joint representations across modalities. We apply self-attention without incorporating cross-modal attention in individual encodings, which reduces computational overhead, fits into our relatively small sample size, and enables a flexible composition of modality inputs. To address the high number of modalities that fall under the transcriptomics upon perturbation category, we insert bottleneck tokens and restrict attention thereafter (Supplementary Fig. S1b). The output bottleneck token embeddings are max-pooled to generate a multimodal drug embedding. Unless explicitly mentioned (as in the case of Madrigal-LLM), a bilinear decoder is used for scoring the probabilities of a pair of compounds having any safety outcomes for computational efficiency.
Denote a flexible fusion module as $g^{\mathrm{fusion}}$, which maps any one or combination of modality-specific encodings for compound $d$: $\{f^{m}(x^{m})\}_{m\in\mathcal{M}_{d}}$ to a single compound encoding $\mathbf{z}^{\mathcal{M}_{d}}\in\mathbb{R}^{\mathrm{joint}}$. Denote as $h^{\mathrm{dec}}:\mathbb{R}^{\mathrm{joint}}\times\mathbb{R}^{\mathrm{joint}}\times\mathcal{R}\to[0,1]$ the decoder for the prediction of safety outcomes from the multimodal encodings of drugs. Let $S$ denote all samples, and $S_{\mathrm{neg}}$ denote all negative samples. We then jointly optimize all encoders $f^{m}$ (initialized from stage 2) and $g^{\mathrm{fusion}},h^{\mathrm{dec}}$ (both randomly initialized), s.t. the loss
$(\displaystyle L_{\mathsf{BCE}}^{\mathrm{ft}}=\frac{1}{|S|}\Bigg{(}$$-\sum_{(d_{1},d_{2},r)\in S}\log h^{\mathrm{dec}}\big{(}g^{\mathrm{fusion}}(\{f^{m}(d_{1}^{m})\}_{m\in\mathcal{M}_{d_{1}}}),g^{\mathrm{fusion}}(\{f^{m}(d_{2}^{m})\}_{m\in\mathcal{M}_{d_{2}}}),r\big{)}$
@@
is minimized.
@@ 2.2.5 Implementation details
At the third stage of model training, we finetuned the encoders, fusion module, and decoder on the combination safety prediction task, with encoders initialized from a pretrained model checkpoint. We used the AdamW optimizer for all three stages and followed a linear warm-up with a cosine annealing schedule, a common practice in training multimodal models.
Hyperparameter tuning. We leverage Weights and Biases [148] to select optimal hyperparameters via a random search over the hyperparameter space. The best-performing hyperparameters are selected by optimizing the AUPRC on the validation set. The hyperparameter space on which we perform a random search to choose the optimal set of hyperparameters is: position embedding $\in[\text{learnable},\text{sinusoidal}]$, position embedding dropout $\in[0.1,0.2,0.4]$, number of heads in transformer encoder $\in[2,4,8]$, dimension of heads in transformer encoder $\in[64,128,256]$, number of layers in transformer encoder $\in[2,3,4,6]$, dimension of feed forward layer in transformer encoder $\in[256,512,1024]$, dropout in transformer encoder $\in[0.2,0.3,0.4]$, number of attention bottlenecks $\in[2,4]$, dropout in projector $\in[0.1,0.2,0.4]$, warmup epochs $\in[10,20,50,100]$, learning rate (for each of structure encoder, pathways encoder, cell viability and transcriptomics encoder, fusion module, decoder) $\in[1e$-$4,5e$-$4,1e$-$3,5e$-3]3]3 ], weight decay $\in[0.001,0.01,0.1]$, epsilon $\in[1e$-$8,1e$-$7,1e$-6]6]6 ], whether or not having a separate adaptor when the drug only has one (structure) modality available, and the ordering of dropout and normalization layer (i.e., normalization before dropout, or vice versa).
@@
[…]
@@
Madrigal is computationally efficient due to the relatively small computation overhead and the feasibility of fitting the entire dataset to a single GPU with efficient operations, taking only a few hours to fine-tune the DrugBank data set on an A100 GPU, compared to the days required by many strong baselines (Supplementary Fig. S2).

### v2

@@ Discussion
Figure 6: Madrigal predicts personalized drug combination efficacy and safety in ex vivo cancer models and real-world patients. a, Using Madrigal to predict individualized drug combination efficacy in ex vivo cancer models. b, Performance of Madrigal in predicting synergistic drug combinations in BeatAML [39], where prediction target is combination synergy (Methods Sec. 4.6). Model evaluation is conducted on randomly held-out patients. Patient-centric and drug-centric denote two ways of calculating AUROC to evaluate the model (Methods Sec. 4.6). Error bars show 95% confidence interval. c, Performance of Madrigal in predicting drug combination efficacy in PDX Encyclopedia[41] when leaving each drug combination out. The prediction target is treatment response (BestAvgResponse). d, Predicted progression-free survival (PFS, TimeToDouble) for individual patient models treated with the (encorafenib + binimetinib) combination. The predictor is trained on other drug combinations, with PFS as the prediction target. Predictions are color-coded by the observed best response category (calculated from response according to mRECIST [41]) of each patient model. PD, progressive disease; SD, stable disease; PR, partial response; CR, complete response. e, Kaplan-Meier survival estimates stratified by predicted treatment response for the (BKM120 + encorafenib) combination. The predictor is trained on other drug combinations with treatment response as the prediction target (same as (d)). f, Same as (e) but for the (encorafenib + binimetinib) combination. g, Using Madrigal to predict personalized drug combination safety in real-world cohorts. h, Model training and inference for predicting drug combination safety in patients. i, Performance of TransformerEHR and TransformerEHR with Madrigal drug embeddings across re-admission prediction, mortality prediction, and adverse event prediction (anemia, hyperglycemia, hyperkalemia, hyponatremia, thrombocytopenia) tasks in the longitudinal event-time cohort. Error bars show standard deviation. j, Performance of combining Madrigal with patient information to predict adverse events for individual patients in the single-index oncology cohort, compared with using Morgan fingerprint or one-hot regimen encoding. Error bars show standard deviation.
@@ 1.1 Drug Combination Safety Dataset
We collect datasets from TWOSIDES (2019-11) [46] and DrugBank (2023-01-04) [47]. TWOSIDES is a database derived from the FDA Adverse Event Reporting System (FAERS). FAERS is a comprehensive repository of adverse event and medication error reports submitted to the FDA. To ensure the reliability and relevance of our data, we adhere to widely accepted criteria in the existing literature on drug safety data mining [105, 106]. These criteria include: (1) a minimum of three reports for the pair of drugs that report the side effect; (2) a proportional reporting ratio of at least 2; (3) a mean reporting frequency of 0.01 or higher; and (4) a Chi-square test statistic of 3.841 or higher ($p$-value <0.05<0.05, with propensity-score matched drugs). Applying these criteria and filtering out safety outcomes with less than 100 samples [50], we have compiled a total of 4,656,138 samples, which include 1,457 unique drugs and 795 unique safety outcomes.
@@ 1.7 MASH Combination Therapies
1.8 Drug Combination Clinical Trials Dataset
To allow for external validation on comprehensive head-to-head drug combination safety comparisons, we systematically extract clinical trials concurrently investigating more than one combination. We extract AE data from clinical trials registered on clinicaltrials.gov through the Continuous Drug Combination Database (CDCDB) [65] (originally sourced via AACT; version April 16, 2024). Two trials not in CDCDB were additionally identified and included through intensive Deep Research using OpenAI’s GPT-o3 model. To ensure valid comparisons of safety, we apply the following criteria for selecting appropriate trials:
  • Is above Phase I;
  • Started after year 2000;
  • Has AE data;
  • Has at least 20 participants per arm on average;
  • Has at least two arms with exactly two different small molecule drugs that can be mapped to our DrugBank identifiers; each arm has a safety population size of at least 20.
We next compare AE data between each arm pair (pairs of arms with exactly two different small molecule drugs and a safety population size of at least 20). We apply three criteria to obtain the data for comparison for each trial (arm pairs and AEs):
  • The AE incidences significantly differ between arms with a two-sided Fisher’s exact test after Bonferroni correction (adjusted $p$-value << 0.05);
  • At least one of the two arms had $\geq$3 affected participants with an incidence $\geq$1%;
  • The two arms should have comparable patient backgrounds and no design confounders such as crossover.
Considering these three criteria, we identify 35 trials (Supplementary Table S10). For trials that contain several arms administering the same drug combination (e.g., at different institutions), we manually confirm that their safety trends were concordant. We thus deem the comparison between that combination and any other combination arm significant for an AE if at least one of the same-combination arms meet the statistical threshold.
@@ 1.10 Patient-derived Xenograft Drug Combination Dataset
The efficacy measures we use are TimeToDouble, which corresponds to progression-free survival (time until tumor volume reaches 200% of baseline), and BestAvgResponse, which corresponds to response (minimum value of the average of $\Delta{Vol}_{t}$ from t == 0 to T, for T $\geq$ 10 d). The drugs in the dataset are mapped to DrugBank ID by name and through manual confirmation with literature.
1.11 Longitudinal Event-Time Cohort
We extract the cohort from EHRSHOT [42], which contains de-identified structured data (e.g., diagnosis and procedure codes, medications, lab values) from EHRs of 6,739 patients from Stanford Medicine. EHRSHOT is longitudinal and includes data beyond ICU and emergency department patients.
We exclude patients with fewer than two visits or those lacking procedure, medication, or diagnosis codes, following [90]. After filtering, 768 patients remain in the cohort for the re-admission and mortality prediction tasks. For the AE prediction task, we use the AE seriousness labels (normal, mild, moderate, severe) provided in [42]. For each patient’s visit and each AE, we select the first occurrence of the highest seriousness (because the patient can be tested multiple times in a visit), resulting in 589, 576, 647, 577, and 612 samples (composed of patient’s visit, time, AE, seriousness) for thrombocytopenia, hyperkalemia, hypoglycemia, hyponatremia, and anemia, respectively.
1.12 Single-Index Oncology Cohort
Analyses of patient-level data from the Dana-Farber Cancer Institute were conducted with approval from the Dana-Farber Institutional Review Board under protocols 19-033 and 19-025. Both protocols were granted waivers of authorization under the Health Insurance Portability and Accountability Act (HIPAA).
We curate a single-index oncology cohort from the Dana-Farber Cancer Institute in which patients were on first-line regimens that contained exactly two small-molecule oncology drugs from June 2015 to March 2025. Regimens are retained only when at least ten patients meet these criteria and neither drugs are indicated for hematological malignancies (as hematological malignancies can confound hematological AE measurements). Patients with missing treatment-related ICD codes or diagnosed with hematological malignancies are excluded. We consider the following AEs: hematotoxicity, neuropathy, thromboembolism, renal impairment, and fluid and electrolyte imbalance. Each AE is mapped to a set of ICD-10 codes determined by an oncology expert (Supplementary Table S15). These ICD-10 codes recorded within 28 days from the start of the first cycle of a regimen are flagged as regimen-related AEs. In total, we curate 3,577 patients with 26 unique regimens and 13 AE types (Supplementary Table S15). For each patient, in addition to regimen and ICD-based tumor tissue type, we also include age, gender, and palliative intent of treatment for each patient.
@@ 2.2.1 Initializing and adapting individual modality-specific encoders
    $L_{\mathsf{BCE}}^{\mathsf{PKG}}=\frac{1}{n_{\mathsf{PKG}}}\Big($ $-\sum_{(s,t,r)\in E}\log h^{\mathsf{PKG}}(f^{\mathsf{PKG}}(s,G),f^{\mathsf{PKG}}(t,G),r)$
@@ 2.2.2 Modality alignment with multimodal contrastive learning
In this stage, our objective is to align the representations generated for the same drug from different modalities with the structure modality with a multimodal contrastive representation learning objective. We adopted the InfoNCE objective [123] with minor modifications similar as in [124], and jointly learn all encoders $f^{m},m\in\mathcal{M}$, initialized from stage 1, s.t. the loss L_cont = ∑_m ≠struc. L(m, struc.) = ∑_m ≠struc. (ℓ(m, struc.) + ℓ(struc., m)), where ℓ(m_u, m_v)=-∑_i=1^B logsim(mu, mv)(di, di)∑j=1B(sim(mu, mv)(di, dj) + 1j ≠i⋅sim(mu, mu)(di, dj) ), and sim_(m_u, m_v)(d_i, d_j) = exp(f^m_u(d_i) ⋅f^m_v(d_j) / τ), is minimized.
@@ 2.2.3 Additional note on modality alignment
[…]
@@ 2.2.4 Model finetuning
Given the impressive performance of attention-based fusion in other multimodal learning contexts, particularly in vision-language models [126, 127, 45], and their flexibility of inputs, we adopt a specialized Transformer encoder architecture with attention bottlenecks for modality fusion to model the joint representations across modalities [43]. To address the large number of cell lines within the transcriptomics modality, we insert bottleneck tokens and restrict attention among those cell line tokens to only within themselves and with the bottleneck tokens thereafter, and vice versa for other modality tokens (Supplementary Fig. S1b). The output bottleneck tokens are max-pooled to generate a multimodal drug embedding. Unless explicitly mentioned (as in the case of Madrigal-LLM), a bilinear decoder is used as the prediction module for scoring the probabilities of a pair of compounds having any safety outcomes for computational efficiency (see Supplementary Note S3 for details).
Denote a flexible fusion module as $g^{\mathrm{fusion}}$, which maps any one or combination of modality-specific encodings for compound $d$: $\{f^{m}(x^{m})\}_{m\in\mathcal{M}_{d}}$ to a single compound embedding $\mathbf{z}^{\mathcal{M}_{d}}\in\mathbb{R}^{\mathrm{joint}}$. Denote as $h^{\mathrm{dec}}:\mathbb{R}^{\mathrm{joint}}\times\mathbb{R}^{\mathrm{joint}}\times\mathcal{R}\to[0,1]$ the prediction module (decoder) for safety outcomes from the multimodal encodings of drugs. Let $S$ denote all samples, and $S_{\mathrm{neg}}$ denote all negative samples. We then jointly optimize all encoders $f^{m}$ (initialized from stage 2) and $g^{\mathrm{fusion}},h^{\mathrm{dec}}$ (both randomly initialized), s.t. the loss
$L_{\mathsf{BCE}}^{\mathrm{ft}}=\frac{1}{|S|}\Bigg($ $-\sum_{(d_{1},d_{2},r)\in S}\log h^{\mathrm{dec}}\big(g^{\mathrm{fusion}}(\{f^{m}(d_{1}^{m})\}_{m\in\mathcal{M}_{d_{1}}}),g^{\mathrm{fusion}}(\{f^{m}(d_{2}^{m})\}_{m\in\mathcal{M}_{d_{2}}}),r\big)$
@@
is minimized. During finetuning, we also randomly drop each available modality with probability 0.5, while ensuring at least one remains. This is equivalent to uniformly sampling one non-empty subset of modalities observed for that compound and further teaches the model to handle whichever modalities available at inference time.
@@ 2.2.5 Implementation details
At the third stage of model training, we finetuned the encoders, fusion module, and prediction module on the combination safety prediction task, with encoders initialized from a pretrained model checkpoint. We used the AdamW optimizer for all three stages and followed a linear warm-up with a cosine annealing schedule, a common practice in training multimodal models. Model checkpoint that achieved the highest AUPRC on the validation set was kept.
Hyperparameter tuning. We leverage Weights and Biases [128] to select optimal hyperparameters via a random search over the hyperparameter space. The best-performing hyperparameters are selected by optimizing the AUPRC on the validation set. The hyperparameter space on which we perform a random search to choose the optimal set of hyperparameters is: position embedding $\in[\text{learnable},\text{sinusoidal}]$, position embedding dropout $\in[0.1,0.2,0.4]$, number of heads in transformer encoder $\in[2,4,8]$, dimension of heads in transformer encoder $\in[64,128,256]$, number of layers in transformer encoder $\in[2,3,4,6]$, dimension of feed forward layer in transformer encoder $\in[256,512,1024]$, dropout in transformer encoder $\in[0.2,0.3,0.4]$, number of attention bottlenecks $\in[2,4]$, dropout in projector $\in[0.1,0.2,0.4]$, warmup epochs $\in[10,20,50,100]$, learning rate (for each of structure encoder, pathways encoder, cell viability and transcriptomics encoder, fusion module, prediction module) $\in[1e$-4,5e4,5e-4,1e4,1e-3,5e3,5e-$3]$, weight decay $\in[0.001,0.01,0.1]$, epsilon $\in[1e$-8,1e8,1e-7,1e7,1e-$6]$, whether or not having a separate adaptor when the drug only has one (structure) modality available, and the ordering of dropout and normalization layer (i.e., normalization before dropout, or vice versa).
@@
  • DrugBank dataset: position embedding $\texttt{sinusoidal}$, position embedding dropout =0.2=0.2, number of heads in transformer encoder =8=8, dimension of heads in transformer encoder =64=64, number of layers in transformer encoder =2=2, dimension of feed-forward layer in transformer encoder =256=256, dropout in transformer encoder =0.3=0.3, number of attention bottlenecks =4=4, dropout in projector =0.1=0.1, warmup epochs =100=100, learning rate (for each of structure encoder, pathways encoder, cell viability and transcriptomics encoder, fusion module, prediction module) =1e=1e-4,1e4,1e-3,1e3,1e-4,1e4,1e-3,1e3,1e-33, weight decay =0.001=0.001, epsilon =1e=1e-66, no separate projector for when the drug has only one (structure) modality available, and normalization layer first.
  • TWOSIDES dataset: position embedding $\texttt{sinusoidal}$, position embedding dropout =0.2=0.2, number of heads in transformer encoder =8=8, dimension of heads in transformer encoder =256=256, number of layers in transformer encoder =2=2, dimension of feed-forward layer in transformer encoder =1024=1024, dropout in transformer encoder =0.2=0.2, number of attention bottlenecks =2=2, dropout in projector =0.2=0.2, warmup epochs =100=100, learning rate (for each of structure encoder, pathways encoder, cell viability and transcriptomics encoder, fusion module, prediction module) =5e=5e-3,5e3,5e-3,1e3,1e-4,1e4,1e-4,1e4,1e-44, weight decay =0.1=0.1, epsilon =1e=1e-77, separate projector for when the drug has only one (structure) modality available, and normalization layer first.
@@
Madrigal is computationally efficient due to the relatively small computation overhead and the feasibility of fitting the entire dataset to a single GPU with efficient operations, taking only a few hours to fine-tune on the DrugBank data set on one GPU, compared to the longer runtime required by a few strong baselines (Supplementary Table S7).